Please help me solve this by multiplying, then simplifying.
(3-2sqrt3)^2

Answer by stanbon(75887)

(3-2sqrt3)^2
(3-2sqrt(3))(3-2sqrt(3))
Use FOIL:
F: 3*3 = 9
O: 3*-2sqrt(3) = -6sqrt(3)
I: 3*-2sqrt(3) = -6sqrt(3)
L: -2sqrt(3)*-2sqrt(3) = 12
--------------------------------
Combine the terms to get
= 9+12 -12sqrt(3)
= 21 -12sqrt(3)
=================
Cheers,
Stan H.